# Copy-Transform-Paste: Zero-Shot Object-Object Alignment Guided by Vision-Language and Geometric Constraints

### v1

@@ 1 Introduction
At test time, we perform a phased optimization that progressively increases the weights of the fractional soft-ICP and penetration terms while zooming cameras toward the interaction region; multiple random restarts with best-of-$N$ selection improve robustness to initialization. We validate on a curated benchmark of 50 mesh-prompt pairs against geometric and LLM-based baselines, observing higher semantic agreement and lower intersection volume.
@@ 3.2 Text Guidance
Let $M_{S}$ and $M_{T}$ be the source and target meshes, respectively. At each step, we update the pose parameters $\theta$ of $M_{S}$ and denote the posed mesh by $\tilde{M}_{S}$. The scene is $S=M_{T}\cup\tilde{M}_{S}$. We render $N$ views $\{I_{i}\}_{i=1}^{N}$ with the differentiable renderer $\mathcal{R}$ from cameras $\{c_{i}\}$ and encode each view and the prompt with CLIP, $e_{i}=\texttt{CLIP}_{\text{img}}(I_{i})$ and $e_{t}=\texttt{CLIP}_{\text{text}}(t)$, and define the text-guidance loss
@@ 3.4 Optimization and Scheduling
Large scale disparities (e.g., a very small source mesh within a large target scene) can dilute gradients from language-vision scoring. Cameras are therefore scheduled across phases to progressively focus on the source mesh. Let $\mathbf{c}_{t}$ be the target-mesh centroid and $\mathbf{c}_{s}^{(p)}$ the source-mesh centroid at phase $p$. The per-phase look-at target interpolates from the target to the source center,
@@ 3.5 Random restarts and per-step noise
Pose optimization is local and sensitive to initialization: if the source mesh starts near an unintended region of the target mesh, the optimizer can converge to a spurious contact. To mitigate this, the procedure launch $N$ independent initializations and runs the same objective for a fixed number of optimization steps. The final estimate is chosen by the highest evaluation score. Fig. 6a illustrates the impact of initialization. Starting from two different random placements of the carrot relative to Bugs Bunny, the optimization converges to distinct yet plausible attachment points. This sensitivity motivates running multiple restarts and selecting the best result under the objective.
@@ 4.1 Experimental Setup
We run 2,0002{,}000 optimization steps per pair with a batch of $8$ camera views per step. Optimization proceeds in P=3P{=}3 phases; across phases, the fractional soft-ICP and penetration weights follow a logarithmic increase, encouraging broad exploration early and stronger contact enforcement later. Camera scheduling (progressive zoom-in and look-at interpolation) follows Sec. 3.4.

### v2

@@ 1 Introduction
At test time, we perform a phased optimization that progressively increases the weights of the fractional soft-ICP and penetration terms while zooming cameras toward the interaction region; multiple random restarts with best-of-NN selection improve robustness to initialization. We validate on a curated benchmark of 50 mesh-prompt pairs against geometric and LLM-based baselines, observing higher semantic agreement and lower intersection volume.
@@ 3.2 Text Guidance
Let $M_{S}$ and $M_{T}$ be the source and target meshes, respectively. At each step, we update the pose parameters $\theta$ of $M_{S}$ and denote the posed mesh by $\tilde{M}_{S}$. The scene is $S=M_{T}\cup\tilde{M}_{S}$. We render NN views $\{I_{i}\}_{i=1}^{N}$ with the differentiable renderer $\mathcal{R}$ from cameras $\{c_{i}\}$ and encode each view and the prompt with CLIP, $e_{i}=\texttt{CLIP}_{\text{img}}(I_{i})$ and $e_{t}=\texttt{CLIP}_{\text{text}}(t)$, and define the text-guidance loss
@@ 3.4 Optimization and Scheduling
Large scale disparities (e.g., a very small source mesh within a large target scene) can dilute gradients from language-vision scoring. Cameras are therefore scheduled across phases to progressively focus on the source mesh. Let $\mathbf{c}_{t}$ be the target-mesh centroid and $\mathbf{c}_{s}^{(p)}$ the source-mesh centroid at phase pp. The per-phase look-at target interpolates from the target to the source center,
@@ 3.5 Random restarts and per-step noise
Pose optimization is local and sensitive to initialization: if the source mesh starts near an unintended region of the target mesh, the optimizer can converge to a spurious contact. To mitigate this, the procedure launch NN independent initializations and runs the same objective for a fixed number of optimization steps. The final estimate is chosen by the highest evaluation score. Fig. 6a illustrates the impact of initialization. Starting from two different random placements of the carrot relative to Bugs Bunny, the optimization converges to distinct yet plausible attachment points. This sensitivity motivates running multiple restarts and selecting the best result under the objective.
@@ 4.1 Experimental Setup
We run 2,0002{,}000 optimization steps per pair with a batch of 88 camera views per step. Optimization proceeds in P=3P{=}3 phases; across phases, the fractional soft-ICP and penetration weights follow a logarithmic increase, encouraging broad exploration early and stronger contact enforcement later. Camera scheduling (progressive zoom-in and look-at interpolation) follows Sec. 3.4.